Government survey takers determine that typical family expenditures each month in the year designated as the base year are as follows:

• 25 pizzas, $10 each

• Apartment rent, $600 per month

• Gasoline and car maintenance, $100 per month

•  Phone service (basic service plus 10 long-distance calls), $50 per month


In the year following the base year, the survey takers determine that pizzas have risen to $11 each, apartment rent is $610, gasoline and maintenance costs are $115, and phone service has dropped in price to $40.


Instructions: Enter your responses by rounding the CPI to three decimal places and the rate of inflation to one decimal place.


a. Find the CPI in the subsequent year and the rate of inflation between the base year and the subsequent year.

CPI:    

Rate of inflation:    %

Here are values of the CPI (multiplied by 100) for each year from 2000 to 2010. For each year beginning with 2001, calculate the rate of inflation from the previous year.

Instructions: Enter your responses rounded to two decimal places.
  

Year CPI Inflation Rate (%)
2000 127.7
2001 133.2
2002 137.3
2003 141.5
2004 145.2
2005 149.4
2006 153.9
2007 157.5
2008 160
2009 163.6
2010 169.2

Compared to 2000–2001, inflation rates remained   (Click to select)   the same   lower   higher  throughout the 2000s.
Answer #1
Year CPI INFLATION RATE (%)
2000 127.7 ------
2001 133.2 [(133.2-127.7) / 127.7 ] * 100 = 4.31
2002 137.3 [(137.3 - 133.2) / 133.2] *100= 3.08
2003 141.5 [(141.5 - 137.3) / 137.3]*100 = 3.06
2004 145.2 [(145.2 - 141.5) / 141.5] *100= 2.61
2005 149.4 [(149.4 - 145.2) / 145.2]*100= 2,89
2006 153.9 [(153.9 - 149.4) / 149.4]*100= 3.01
2007 157.5 [(157.5 - 153.9)/153.9]*100= 2.34
2008 160 [(160 - 157.5)/157.5]*100= 1.59
2009 163.6 [(163.6 - 160) /160]*100= 2.25
2010 169.2 [(169.2 - 163.6) / 163.6]*100= 3.42

Inflation rate =[( current year CPI - Previous year CPI ) / Previous year CPI] * 100

Compared to 2000-2001, inflation rates remained lower throughout the 2000s.